# Cosmological perturbations of a relativistic MOND theory

### v1

@@ V Perturbation theory
where $\sigma$ is a function of spacetime with arbitrary amplitude; in PN expansion $\Sigma$ and other higher order expansions in Eq. (23) are absorbed to $\sigma$; if needed we use an overbar to indicate the background order quantity. The fully nonlinear and exact perturbation formulation includes the PN expansion; comparing Eqs. (62) and (21) we have
@@
with ${\cal N}$ given in Eq. (95). Apparently, the expansion is lengthy if we keep perturbations of $\sigma$. In the following, we consider $\sigma$ in the linear order perturbation theory, and for simplicity, set $\sigma=0$ in the fully nonlinear and exact formulation; $\sigma=0$ corresponds to a temporal gauge condition valid to fully nonlinear order and it is the same as the comoving gauge of the $\tau$-fluid, see Eqs. (72) and (99).
@@ V.1 Linear perturbation
where $\alpha_{\sigma}$ is a unique gauge-invariant notation which becomes $\alpha$ in the $\sigma=0$ gauge condition, see later. The $\tau$-field energy-momentum tensor becomes
@@
To the background order, we recover Eq. (13), and only scalar-type perturbations are excited without anisotropic stress. We have
${\delta p\over\delta\mu}=c_{\tau}^{2}\Big{(}1-{2a^{3}c_{\tau}^{2}\over I_{0}\delta Q}{\cal J}_{,A}{\Delta\over a^{2}}\Upsilon\Big{)}^{-1}.$ (73)
The entropic perturbation of $\tau$-fluid is
@@
$\hskip-22.76228pt\dot{\kappa}+2H\kappa+\Big{(}c^{2}{\Delta\over a^{2}}+3\dot{H}\Big{)}\alpha=-{4\pi G\over 3c^{2}}(\delta\mu+\delta p),$ (78)
@@
where the fluid quantities are the collective ones, for example, $\delta\mu=\delta\mu_{m}+\delta\mu_{\tau}$, $\delta p=\delta p_{m}+\delta p_{\tau}$, $(\mu+p)v=(\mu_{m}+p_{m})v_{m}+(\mu_{\tau}+p_{\tau})v_{\tau}$; we decomposed $\chi_{i}\equiv\chi_{,i}+\chi_{i}^{(v)}$ with $\chi^{(v)i}_{\;\;\;\;\;\;,i}\equiv 0$. We ignored the anisotropic stress of ordinary fluid in Eq. (79). The vector- and tensor-type perturbation equations are not affected by the $\tau$-field.
@@
These two equations are valid for ordinary fluid with the anisotropic stress ignored in the second equation, and are also valid for the $\tau$-fluid with the $\tau$-fluid quantities given in Eq. (72). Using the $\tau$-fluid quantities, Eq. (81) is naturally valid by the equation of motion in Eq. (80) together with $\dot{\cal J}=0$. Using Eqs. (16), (66) and (72), Eq. (82) is identically valid.
The complete set of linear perturbation equations without imposing temporal gauge condition is Eqs. (75)-(79) from Einstein’s equation, Eqs. (81) and (82) for other fluids, and the same equation for the $\tau$-fluid with the equation of state in Eqs. (72) and (74). For the $\tau$-fluid, instead of conservation equations in (81) and (82), we can directly use the field variable and the equation of motion in Eq. (80).
@@ V.1.1 Gauge transformation
As a temporal gauge condition we can use $\sigma=0$ in all coordinate systems, and we may call it uniform-$\tau$ gauge. This implies $v=0$ for the $\tau$-fluid, and we may call it the comoving gauge of the $\tau$-fluid. Similarly we can impose $v=0$ for an ordinary fluid. As another choice we can set $\delta\mu=0$ in all coordinates which is the uniform-density gauge for the $\tau$-fluid or ordinary fluid. Other temporal gauge conditions which remove the gauge degree of freedom completely are the following: $\varphi\equiv 0$ (uniform-curvature gauge), $\kappa\equiv 0$ (uniform expansion gauge), $\chi\equiv 0$ (zero-shear gauge) Bardeen-1988 . One exception where the temporal gauge degree of freedom is not completely fixed even after imposing the gauge condition is $\alpha\equiv 0$ (synchronous gauge). These gauge conditions are applicable to nonlinear and exact order with corresponding gauge invariance for other than synchronous gauge Noh-Hwang-2004 ; Hwang-Noh-2013 .
@@ V.1.2 Density perturbation in $\sigma=0$ gauge
The ${\cal K}(Q)$-term is introduced in the Lagrangian in order to reproduce cosmology which demands presence of dark matter in the context of CMB and the large-scale structure. Here, we are curious whether the pressure terms in the above equation can be ignored in the perturbation level so that the $\tau$-field in the form of ${\cal K}$-term can supplement the large-scale dark matter demanded in current cosmology.
@@
where the right-hand side is evaluated in the background; we used Eqs. (16), (17), (72), (66) and (67). The density perturbation of $\tau$-field is stable (unstable) for $k>k_{J}$ ($k<k_{J}$), and it behaves as a cold dark matter for negligible presure perturbation with $k\ll k_{J}$. The above Jeans criterion is valid for general ${\cal J}$ and ${\cal K}$, assuming the background $\tau$-field behaves as a dust fluid.
@@ V.1.3 Perturbation constraint
$\delta\mu_{\sigma}=-{c^{4}\over 8\pi G}\Big{\{}n{I_{0}\over a^{3}}\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{-1/n}$
×$\qquad\times\Big{[}1+\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n}\Big{]}^{2}\alpha+2{\cal J}_{,A}{\Delta\over a^{2}}\Upsilon\Big{\}},$
$\delta p_{\sigma}=-{c^{4}\over 8\pi G}{I_{0}\over a^{3}}\Big{[}1+\Big{(}{I_{0}\over 2\nu^{2}{\cal K}_{n+1}a^{3}}\Big{)}^{1/n}\Big{]}\alpha.$ (90)
For a dust-like equation of state for the background, we have $[I_{0}/(\nu^{2}{\cal K}_{n+1}a^{3})]^{1/n}(\ll 1)$ and Eq. (88) is valid.
The Jeans scale becomes
@@
thus, the Jeans scale corresponds to the length scale of the mass-like term. In Eq. (48) a reasonable phenomenology in galactic scale demanded $1/\nu>1{\rm Mpc}$ and this applies only for $n=1$. In Eq. (20) a dust-like behavior in the background demands $1/\nu\leq 0.22kpc$. Thus, $n=1$ case is excluded Blanchet-Skordis-2024 .
@@
$L_{J}\leq\sqrt{{L_{H}L_{*}\over\sqrt{6\Omega_{\tau 0}a}}(1+{\cal J}_{,A})}\sim\sqrt{{1+{\cal J}_{,A}\over\sqrt{a}}}1.1Mpc,$ (94)
where we used $k_{J}\sim L_{J}^{-1}$, $L_{H}\equiv c/H_{0}=4.3Gpc$ and Eq. (20). In non-MOND regime ${\cal J}_{,A}$ should be negligible, see Eq. (46) and below.
@@ V.2 Fully nonlinear and exact perturbations in $\sigma=0$ gauge
$U_{0}=-a{\cal N},\quad U_{i}=0,$
$U^{0}={1\over a{\cal N}},\quad U^{i}={\chi^{i}\over a^{2}{\cal N}(1+2\varphi)}.$ (96)
@@
For fully nonlinear and exact formulation including the transverse (vector) and transverse-tracefree (tensor) modes to fully nonlinear and exact order, see Gong-2017 . The fluid quantities in Einstein’s equations are the collective ones, for example, $E=E_{m}+E_{\tau}$, $m_{i}=m_{i}^{m}+m_{i}^{\tau}$, $m_{ij}=m_{ij}^{m}+m_{ij}^{\tau}$, etc; for the $\tau$-fluid, fluid quantities are given in Eq. (99), and for ordinary fluid, scalar field, and Maxwell field, see Sec. IV.C in Hwang-Noh-EM-2023 .
@@
The conservation equations are valid for $\tau$-fluid and ordinary fluid, including scalar field as well as Maxwell field Hwang-Noh-EM-2023 . The equation of motion in Eq. (7) gives
@@
Ignoring the gravitational waves, Eqs. (100)-(108) are the complete set of exact perturbation equations. The conservation equations in (106)-(107) are valid for fluids as well as for $\tau$-field with the equation of state presented in Eq. (99); for the field we can also use the equation of motion in Eq. (108).
@@ VI Discussion
The PN approximation of the BMS theory was presented in Flanagan-2023 ; Blanchet-Skordis-2024 in Minkowsky background, and the cosmological linear perturbation was presented in Blanchet-Skordis-2024 . Compared with previous works, here we presented equations valid to 1PN order in cosmology context, linear perturbations without imposing temporal gauge condition, and the fully-nonlinear and exact perturbation equations. We clarified the gauge issues in both formulations. We presented the 1PN approximation assuming $\tau=t$ as a physical condition and fully nonlinear and exact perturbations in $\tau=\tau(t)$ using a temporal gauge condition. General cases without imposing these conditions demand lengthy presentations, but the methods are the same.
@@
In linear perturbations, we kept general ${\cal J}$ and ${\cal K}$ and derived Jeans criterion for a constraint on $\tau$-fluid. The Bekenstein and Milgrom formulation of MOND Bekenstein-Milgrom-1984 can be recovered by using ${\cal J}(A)$ in the 0PN limit of the PN approximation by taking $\sigma=0$ as a physical condition on $\tau$-field, thus $\tau=t$. The result is derived without assuming stationary condition mentioned in Flanagan-2023 ; Blanchet-Skordis-2024 .
@@
Constraints on ${\cal K}(Q)$ follow from the background, perturbations, and the MOND requirements. (i) A demand for dust-like equation of state for $\tau$-field in the background constrains ${\cal K}$: see Eq. (20) for model-I. (ii) For model-II, a mass-like term appears for $n=1$, and the upper limit on this term gives a constraint in Eq. (48) which conflicts with the background constraint. These two constraints were presented in Blanchet-Skordis-2024 . (iii) A demand for $\tau$-fluid to behave as a pressureless fluid in linear density perturbation leads to a constraint. The Jeans scale is derived in Eq. (89) for general ${\cal J}$ and ${\cal K}$, and in Eq. (91) for model-I. Proper estimation of the effect of ${\cal K}$ in cosmological linear perturbations needs numerical study of the full system summarized below Sec. V.1.
@@
In BMS theory, the two functions ${\cal J}$ and ${\cal K}$ are still free to accommodate the phenomena in MOND and cosmology. Future studies may design and constrain the functional forms using theories and observations. Other functional forms, especially ${\cal K}$ for cosmology were previously considered with successful fits to the power spectra of the large-scale matter and CMB Blanchet-Skordis-2024 ; Skordis-Zlosnik-2021 . The fully nonlinear and exact perturbation equations will be useful to study the leading nonlinear matter power spectrum as well as the non-Gaussian effects.
@@
If successful, by adding two functions of a time-like vector field in a gradient form, the BMS theory encompasses the MOND phenomenon within Einstein’s gravity paradigm algebraically, although the local Lorentz invariance and the strong equivalence principle are broken conceptually. The latter differences may open new possibilities to test the theory.

### v2

@@ V Perturbation theory
where $\sigma$ is a function of spacetime with arbitrary amplitude; in PN expansion $\Sigma$ and other higher order expansions in Eq. (23) are absorbed to $\sigma$; if needed we use an overbar to indicate the background order quantity. The fully nonlinear and exact perturbation formulation includes the PN expansion; comparing Eqs. (69) and (21), we have
@@
with ${\cal N}$ given in Eq. (105). Apparently, the expansion is lengthy if we keep perturbations of $\sigma$. In the following, we consider $\sigma$ in the linear order perturbation theory, and for simplicity, set $\sigma=0$ in the fully nonlinear and exact formulation; $\sigma=0$ corresponds to a temporal gauge condition valid to fully nonlinear order and is the same as the comoving gauge of the $\tau$-fluid, see Eqs. (79) and (109).
@@ V.1 Linear perturbation
where $\alpha_{\sigma}$ is a unique gauge-invariant notation which becomes $\alpha$ in the $\sigma=0$ gauge condition, see later. To the linear order, we have $A=0$. However, in the MOND regime, from Eq. (46), we need ${\cal J}_{,A}=-1+(ac^{2}/a_{M})\sqrt{A}$ and $\sqrt{A}$ is linear order. We have $\sqrt{A}=(\Upsilon^{,i}\Upsilon_{,i})^{1/2}/a=|\nabla\Upsilon|/a$, thus ${\cal J}_{,A}=-1+|\nabla\Upsilon|c^{2}/a_{M}$ which is consistent with Eq. (44) in the MOND regime by identifying $\Upsilon=\alpha_{\sigma}=\Phi/c^{2}$. In Newtonian regime, we have ${\cal J}_{,A}=0$.
The $\tau$-field energy-momentum tensor becomes
@@
To the background order, we recover Eq. (13). The perturbed parts are
$\delta\mu={c^{4}\over 8\pi G}\Big{(}Q{\cal K}_{,QQ}\delta Q-2{\cal J}_{,A}{\Delta\over a^{2}}\Upsilon\Big{)},$
$\delta p={c^{4}\over 8\pi G}{\cal K}_{,Q}\delta Q,\ \quad v={1\over aQ}\sigma.$ (80)
Only scalar-type perturbations are excited without anisotropic stress. We have
${\delta p\over\delta\mu}=c_{\tau}^{2}\Big{(}1-{2c_{\tau}^{2}\over{\cal K}_{,Q}\delta Q}{\cal J}_{,A}{\Delta\over a^{2}}\Upsilon\Big{)}^{-1}.$ (81)
The entropic perturbation (isotropic stress) of $\tau$-fluid is
@@
$\hskip-22.76228pt\dot{\kappa}+2H\kappa+\Big{(}c^{2}{\Delta\over a^{2}}+3\dot{H}\Big{)}\alpha={4\pi G\over c^{2}}(\delta\mu+3\delta p),$ (86)
@@
where the fluid quantities are the collective ones, for example, $\delta\mu=\delta\mu_{m}+\delta\mu_{\tau}$, $\delta p=\delta p_{m}+\delta p_{\tau}$, $(\mu+p)v=(\mu_{m}+p_{m})v_{m}+(\mu_{\tau}+p_{\tau})v_{\tau}$; the matter part include baryon, photon, neutrino, etc.; we decomposed $\chi_{i}\equiv\chi_{,i}+\chi_{i}^{(v)}$ with $\chi^{(v)i}_{\;\;\;\;\;\;,i}\equiv 0$. We ignored the anisotropic stress of ordinary fluid in Eq. (87). The vector- and tensor-type perturbation equations are not affected by the $\tau$-field.
@@
These two equations are valid for ordinary fluid with the anisotropic stress ignored in the second equation; we have $p=\delta p=0$ for dust and $p={1\over 3}\mu$ and $\delta p={1\over 3}\delta\mu$ for photon fluid, etc. These are also valid for the $\tau$-fluid with the $\tau$-fluid quantities given in Eq. (79). Using the $\tau$-fluid quantities, Eq. (89) is naturally valid by the equation of motion in Eq. (88) together with Einstein’s equation and $\dot{\cal J}=0$. Using Eqs. (16), (73) and (79), Eq. (90) is identically valid.
The complete set of linear perturbation equations without imposing temporal gauge condition is Eqs. (83)-(87) from Einstein’s equation, Eqs. (89) and (90) for other fluids, and the same equation for the $\tau$-fluid with the fluid quantities in Eq. (79). For the $\tau$-fluid, instead of conservation equations in (89) and (90), we can directly use the field variable and the equation of motion in Eq. (88).
@@ V.1.1 Gauge transformation
As a temporal gauge condition we can use $\sigma=0$ in all coordinate systems, and we may as well call it uniform-$\tau$ gauge. This implies $v=0$ for the $\tau$-fluid, and we may as well call it the comoving gauge of the $\tau$-fluid. Similarly we can impose $v=0$ for an ordinary fluid. As another choice we can set $\delta\mu=0$ in all coordinates which is the uniform-density gauge for the $\tau$-fluid or ordinary fluid. Other temporal gauge conditions which remove the gauge degree of freedom completely are the following: $\varphi\equiv 0$ (uniform-curvature gauge), $\kappa\equiv 0$ (uniform expansion gauge), $\chi\equiv 0$ (zero-shear gauge) Bardeen-1988 . One exception where the temporal gauge degree of freedom is not completely fixed even after imposing the gauge condition is $\alpha\equiv 0$ (synchronous gauge). These gauge conditions are applicable to nonlinear and exact order with corresponding gauge invariance for other than synchronous gauge Noh-Hwang-2004 ; Hwang-Noh-2013 .
@@ V.1.2 Density perturbation in $\sigma=0$ gauge
The ${\cal K}(Q)$-term is introduced in the Lagrangian in order to reproduce cosmology which often demands presence of dark matter in the context of CMB and the large-scale structure. Here, we are curious whether the pressure terms in the above equation can be ignored in the perturbation level so that the $\tau$-field in the form of ${\cal K}$-term can supplement the dark matter in the large-scale demanded in standard cosmology.
@@
where the right-hand side is evaluated using the background parameters; we used Eqs. (16), (17), (73), (74) and (79). The pressure perturbation can be ignored in super-Jeans scale, and the density perturbation of $\tau$-field behaves as a cold dark matter for negligible pressure perturbation with $k\ll k_{J}$. The above Jeans criterion is valid for general ${\cal J}$ and ${\cal K}$, assuming the background $\tau$-field behaves as a dust fluid.
@@ V.1.3 Perturbation constraint
$\delta\mu_{\sigma}=-{c^{4}\over 8\pi G}\Big{(}n{I_{0}\over a^{3}}{Q^{2}\over Q-1}+2{\cal J}_{,A}{\Delta\over a^{2}}\Big{)}\alpha,$
$\delta p_{\sigma}=-{c^{4}\over 8\pi G}{I_{0}\over a^{3}}Q\alpha.$ (98)
Equation (97) gives
${k_{J}\over a}=\Big{[}{2a^{3}\over I_{0}}{Q-1\over nQ^{2}}\Big{(}{(n+1)Q\over nQ+1}+{\cal J}_{,A}\Big{)}\Big{]}^{-1/2}.$ (99)
For $Q\simeq 1$, using Eq. (19), we have
${\delta p_{\sigma}\over\mu}=c_{\tau}^{2}\Big{(}1-{\cal J}_{,A}{c_{\tau}^{2}c^{2}k^{2}\over 4\pi G\varrho_{\tau}a^{2}}\Big{)}^{-1}\delta_{\sigma},$
${c_{\tau}^{2}c^{2}k_{J}^{2}\over 4\pi G\varrho_{\tau}a^{2}}(1+{\cal J}_{,A})=1.$ (100)
The Jeans scale can be written as
@@
thus, the Jeans scale corresponds to the length scale of the mass-like term. In Eq. (49) a reasonable phenomenology in galactic scale demanded $1/\nu>1{\rm Mpc}$ and this applies only for $n=1$. In Eq. (20) a dust-like behavior in the background demands $1/\nu\leq 220pc$. Thus, $n=1$ case is excluded Blanchet-Skordis-2024 .
@@
${\lambda_{J}\over 2\pi}={a\over k_{J}}\leq\sqrt{{L_{H}L_{*}a^{3/2}\over\sqrt{6\Omega_{\tau 0}}}(1+{\cal J}_{,A})}$
$\qquad\sim\sqrt{1+{\cal J}_{,A}}a^{3/4}0.46Mpc,$ (104)
where we used $L_{H}\equiv c/H_{0}=4.3Gpc$ and Eq. (20). In Newtonian regime we may set ${\cal J}_{,A}=0$ whereas in the MOND regime we have $1+{\cal J}_{,A}={1\over a}|\nabla\Phi|/a_{M}$. Above the Jeans scale the density perturbation of $\tau$-field behaves as a cold DM.
@@ V.2 Fully nonlinear and exact perturbations in $\sigma=0$ gauge
$U_{0}=-a{\cal N},\quad U_{i}=0,\quad U^{0}={1\over a{\cal N}},\quad U^{i}={\chi^{i}\over a^{2}{\cal N}(1+2\varphi)}.$ (106)
@@
To the linear order Eqs. (110)-(114) gives Eqs. (84)-(87) and (83), respectively. For fully nonlinear and exact formulation including the transverse (vector) and transverse-tracefree (tensor) modes, see Gong-2017 . The fluid quantities in Einstein’s equations are the collective ones, for example, $E=E_{m}+E_{\tau}$, $m_{i}=m_{i}^{m}+m_{i}^{\tau}$, $m_{ij}=m_{ij}^{m}+m_{ij}^{\tau}$, etc; for the $\tau$-fluid, fluid quantities are given in Eq. (109), and for ordinary fluid, scalar field, and Maxwell field, see Sec. IV.C in Hwang-Noh-EM-2023 .
@@
The conservation equations are valid for $\tau$-fluid and ordinary fluids, including scalar field as well as Maxwell field Hwang-Noh-EM-2023 . The equation of motion in Eq. (7) gives
@@
Equations (110)-(118) are the complete set of exact perturbation equations. The conservation equations in (116)-(117) are valid for fluids as well as for $\tau$-field with the equation of state presented in Eq. (109); for the field we can also use the equation of motion in Eq. (118).
@@ VI Discussion
The PN approximation of the BMS theory was presented in Flanagan-2023 ; Blanchet-Skordis-2024 in Minkowsky background, and the cosmological linear perturbation was presented in Blanchet-Skordis-2024 . Compared with previous works, here we presented equations valid to 1PN order in cosmology context, linear perturbations without imposing temporal gauge condition, and the fully-nonlinear and exact perturbation equations. We clarified the gauge issues in both formulations. We presented the 1PN approximation assuming $\tau=t$ as a physical condition and fully nonlinear and exact perturbations assuming $\tau=\tau(t)$ using a temporal gauge condition. General cases without imposing these conditions demand lengthy presentations, but the methods are the same.
@@
In linear perturbations, we kept general ${\cal J}$ and ${\cal K}$ and derived Jeans criterion for a constraint on $\tau$-fluid. The Bekenstein and Milgrom formulation of MOND Bekenstein-Milgrom-1984 can be recovered by using ${\cal J}(A)$ in the 0PN limit of the PN approximation by taking $\sigma=0$ as a physical condition on $\tau$-field, thus $\tau=t$.
@@
Constraints on ${\cal K}(Q)$ follow from the background, perturbations, and the MOND requirements. (i) A demand for dust-like equation of state for $\tau$-field in the background constrains ${\cal K}$: see Eq. (20) for model-I. (ii) For model-II, a mass-like term appears for $n=1$, and the upper limit on this term gives a constraint in Eq. (49) which conflicts with the background constraint. These two constraints were presented in Blanchet-Skordis-2024 . (iii) A demand for $\tau$-fluid to behave as a pressureless fluid in linear density perturbation leads to another constraint. The Jeans scale is derived in Eq. (97) for general ${\cal J}$ and ${\cal K}$, and in Eq. (101) for model-I. Proper estimation of the effect of ${\cal K}$ in cosmological linear perturbations needs numerical study of the full system summarized below Sec. V.1.
@@
In BMS theory, the two functions ${\cal J}$ and ${\cal K}$ are still free to accommodate the phenomena in MOND and successful cosmology. Future studies may design and constrain the functional forms using theories and observations. Other functional forms, especially ${\cal K}$ for cosmology were previously considered with successful fits to the power spectra of the large-scale matter and CMB Blanchet-Skordis-2024 ; Skordis-Zlosnik-2021 . The fully nonlinear and exact perturbation equations will be useful to study the leading nonlinear matter power spectrum as well as the non-Gaussian effects.
Recent JWST observation of the early emergence of massive galaxies in high redshifts is regarded as a challenge to the cold dark matter (CDM) cosmology Carniani-2024 . The early formation of galaxies was a predicted feature in the MOND model Sanders-1998 ; McGaugh-2024 . In the low-acceleration MOND regime the gravity is stronger than Newtonian gravity, thus the structures can grow faster Nusser-2002 . We showed the case in the BMS model using the 0PN approximation in Section IV.1.4.
We point out a couple of caveats used in our PN approximation. Our PN approximation is limited by adopting $\tau=t$ to the background order, thus excluding proper consideration of the $\tau$-field contribution to the Friedmann background. Adopting $\sigma=0$ as a physical condition to 0PN order is another, a sort of, ansatz limiting the potential role of the $\tau$-field. The first condition limits our study valid in the Friedmann background without ${\cal K}$ contribution, see below Eq. (23). The latter condition conveniently leads us to the MOND phenomenology but its meaning should be clarified. These are interesting issues demanding further investigation by relaxing the conditions.
We showed that above Jeans scale the density perturbation of $\tau$-field behaves as a pressureless fluid. To confirm whether the Jeans scale in Eq. (97) [Eq. (101) for the model-I and Eq. (104) for $n=2$] is sufficient to reproduce the large-scale structure and CMB observations which are consistent with the $\Lambda$CDM paradigm, we may need numerical study by including contributions of the baryon, photons and other cosmological paraphernalia.
@@
If successful, by adding two functions of a time-like vector field in a gradient form, the BMS theory encompasses the MOND phenomenon within Einstein’s gravity paradigm algebraically, although the local Lorentz invariance and the strong equivalence principle are broken conceptually. The latter violations may open new possibilities to test the theory.